# references.bib

### v1

@@ 2.1 Conceptualizing the mechanisms for baseline regulatory costs across systems
This result is supported by the observation that in unicellular eukaryotes, the number of regulatory genes grows as $G^{1.26}$ [van2003scaling], which is much closer to one than the scaling observed in bacteria. Unlike bacteria, unicellular eukaryotes have various internal spatial partitions, including the partitioning of genes between the nucleus and mitochondria. The transition from prokaryotes to eukaryotes illustrates how the internal structure of organisms, hierarchy, and partitioning can alter the requirements for regulation.
Also, it should be noted that the scaling of $r_{opt}$ in Equation 10 can not continue indefinitely because eventually $I=0$. This limit is given by
$r_{\text{max}}=\frac{1}{4\theta}\left(\frac{\gamma_{3}^{2}}{\gamma_{1}^{2}\eta^{2}\rho}-\rho\right).$ (12)
The number of regulators, $r$, saturates to a constant and thus should have effectively sublinear scaling over much of the range of size. Similarly, in the transition between only regulators and compartments, we expect $r$ to scale slower than quadratically with $f$, especially if $f_{r}\neq f_{c}$, which occurs for a wide variety of parameter combinations. In addition, it is important to note that the unit costs, $\gamma_{i}$, could vary with attributes of the system. For example, the unit costs of compartments in cells depend on the surface area of these compartments. Similarly, the rate and cost of the negative interaction, $\rho$ and $\gamma{1}$ respectively, may change with system size. In Equation 10 such cost scaling would cause $r_{opt}$ to scale faster or slower than $f$. Taking all of these considerations into account, we should expect real-world systems in the most general sense to follow
@@
where $r_{0}$ is the minimum cost for the smallest size ($N=1$), and $\beta$ is the scaling exponent. The above framework gives us the common expectation of $\beta=2$ for a compartment-free system, and $\beta=1$ for a system with both regulators and compartments. We can interpret observed exponents that are closer to $2$ as being well-mixed and compartment-free. Close to $1$ indicates a system optimized on compartments with a mixed strategy. An exponent significantly larger than $1$ and significantly smaller than $2$ indicates the transition from regulators only to some compartmentalization. Exponents significantly less than $1$ indicate the regime where compartments are expanding and regulators are becoming cumbersome and saturating, or where there are certain economies of scaling in the unit costs.
@@ 2.2 Empirical results
Figure 2: Examples of scaling of regulatory functions in bacteria, cities, and Norwegian companies (a) The number of transcription regulatory genes vs. the number of genes in the genome of bacteria [van2003scaling]. (b) The number of lawyers vs. working population in US Metropolitan Statistical Areas. (c) The number of managers in companies in Norway.
We collect data on regulatory components and system size in both biological and socioeconomic systems to examine how regulatory costs scale with size. In biological systems, we focus on bacterial and unicellular eukaryote cells, specifically collecting data on the number of regulatory genes in their genomes. Regulatory genes are those that determine the timing and environmental conditions for gene expression by producing proteins that bind to other genes. We use data reported by [van2003scaling] for these biological systems. For socioeconomic systems, we gathered data on the number of lawyers in cities and the number of managers in Norwegian and Korean companies, US federal government agencies, and various types of US universities. We also synthesize results from another study measuring the number of administrators for Wikipedia pages [yoon2023makes]. For human systems, system size is measured by the population of the entity, such as the number of employees for companies and universities, and the number of editors for Wikipedia pages. The scaling exponents of these systems are estimated using Eq. 13, with $\beta$ representing the scaling exponent. For detailed information on data sources and statistical methods used for estimation, see Data and Methods.
@@
In the data we gathered, regulatory costs scale with system size across diverse systems. Three examples are shown in Fig. 2: regulatory genes in bacteria, lawyers in cities, and managers in Norwegian companies. As predicted, the scaling exponents vary across system types, with regulatory genes in bacteria having the highest exponent at $1.84$. The number of lawyers scales superlinearly with the urban population, with an exponent of $1.28$ (Fig. 2B). This is similar to prokaryotic regulatory genes but does not reach quadratic scaling due to constraints on individuals’ social interaction capacity and the overall network structure of cities. Theories of urban scaling based on these constraints have successfully predicted similar exponents for many socioeconomic outputs driven by interactions [bettencourt2013origins, yang2019modeling]. Contrary to cities, human organizations—-such as government agencies, companies, and universities—-typically exhibit a high degree of hierarchical structure, leading us to expect different scaling behaviors. In these organizations, managers play a crucial role in coordinating efforts and mitigating conflicts among subordinates. The scaling exponent for managers in Norwegian companies is 0.91 (Fig. 2C).
@@
The scaling exponents for all data we have gathered, including the number of managers in all sectors of US universities, Norwegian and Korean companies, and US federal government agencies, as well as those in biological systems, are summarized in Fig 3. In all the modular systems, the number of managers scales sublinearly with small variations—from $0.94$ for Federal agencies (highest exponent) to $0.72$ for Korean companies (lowest exponent). The sublinear scaling shown in the data suggests the span of control, the number of subordinates per manager, increases with organization size. This finding aligns with previous studies of companies and public agencies in management science [boyne2013burdened], while is contrary to the popular belief that larger organizations are less efficient in coordination costs [williamson1967hierarchical]. Even in Wikipedia, a supposedly decentralized system, the interaction network among editors has been shown to be highly modularized [yoon2023makes]. Interactions tend to cluster around specific topics of interest, allowing a smaller number of administrators to effectively manage issues that arise. As such, despite the decentralized nature of the system, administrators naturally oversee specific modules and thus sublinearly scale with the total contributors.
@@
Figure 3: The scaling exponents ($\beta$) and 95% confidence intervals for regulatory functions in various systems. With more well-mixed internal structures, such as bacteria cells and cities, regulatory functions tend to be superlinear. With more modular internal structures, such as companies and universities, the scaling exponents are linear to sublinear.
@@ 2.3 Function diversity associated with scaling deviations
Figure 4: Industry and Federal agencies colored by occupations’ normalized entropy, a metric for the diversity of functions in the system. Higher values denote more diversity. The inset shows the correlation between the normalized entropy and scaling deviation. Systems that are more diverse have more managers than would be expected of their size. The Pearson correlation is 0.59 for industries and 0.20 for federal agencies.
Motivated by our theory, we predict that greater functional diversity is associated with higher regulatory costs. To test this prediction, we will quantify functional diversity across a range of systems and demonstrate that it is positively associated with the scaling residuals. In other words, systems with higher functional diversity exceed the expected number of regulators according to the scaling curve, while those with lower functional diversity fall short.
We quantify functional diversity in an organization by analyzing the distribution of occupations within it. For this analysis, we utilize individual-level occupation information for US federal government agencies. To ensure robustness, we also test our predictions on companies. Although occupation information is not available at the company level, we study companies aggregated into industries, for which we have detailed accounts of the distribution of occupations. We measure functions using the finest occupational categories available in these datasets. For more details on data sources and occupation definitions, see Data and Methods.
@@
We also compute the scaling residual [bettencourt2010urban] for each system, which quantifies the extent to which a system over- or under-performs relative to the scaling curve. The scaling residual for system $i$, $\xi_{i}$, is defined as, $\xi_{i}=\log r_{i}-\log r(N_{i})$, where $r_{i}$ is the number of regulators for system $i$ in the data, and $r(N_{i})$ is the regulators expected of its size according to Eq. 13.
Figure 4 illustrates the relationship between managers and employees, with each entity colored according to its normalized entropy, where higher values indicate greater function diversity [yang2022scalinguniversalityfunctiondiversity]. For both federal agencies and industries, entities with higher function diversity tend to have more managers than expected for their size. The insets of Fig. 4 display the correlation between scaling residuals and normalized entropy. The correlation is 0.59 for industries ($p<0.001$), and .20 for federal agencies ($p=0.030$). These results support our initial hypothesis that greater function diversity is associated with greater regulatory costs due to greater coordination requirements.
@@ 3 Discussion
We have proposed a conceptual framework for unifying the prediction of regulatory costs across biological and socioeconomic systems by examining the interactions of fundamental features, size and internal structure. We also conducted an empirical analysis of regulatory costs across these systems. Our findings indicate that the variation of regulatory functions with system size depends significantly on the system’s structure. The exponents range from nearly 2 for regulatory genes in bacteria, which have a well-mixed internal structure, to sublinear scaling for managers in hierarchically organized human organizations, such as federal agencies and universities. By characterizing regulatory functions based on their underlying mechanisms rather than the system type, our work represents a first step toward a unified understanding of maintenance functions across diverse systems.
@@
Our conceptual model is based on the simplest assumptions about cost factors, intended to lay the groundwork for more advanced theories in the future, These future models should incorporate more accurate cost estimates tailored to different system types through empirical measurements. For example, in hierarchical human organizations, the cost of regulators within compartments should also be factored into future models. In biological systems, more precise cost calculations should consider the frequency of gene expression, which plays a critical role in determining the energetic cost in the cell.
@@
The greater scaling exponents of biological systems do not necessarily imply biological systems are less efficient than socioeconomic ones. Creating compartmentalization and structure in biology is expensive—intricate physical structures need to be developed to separate the cell nuclei from the rest. Extending this idea, complex physical infrastructure also needs to be developed to enable organs in animals, where genes only get expressed in a certain tissue. While bacteria need to carry a rapidly increasing number of regulatory genes with increasing size, they save on the energetic cost of creating compartmentalization within the cell. In comparison, developing structure in social systems does not necessarily incur a physical cost. For example, a company’s CEO can decide to create a new division in the company without employing new physical separations between divisions, and the re-organization can be accomplished in a matter of weeks. Humans are also naturally creatures of groups with limited social capacity, making modularity a common characteristic in many social systems, even decentralized ones such as cities. Indeed, the social interactions scale with $N^{1.2}$ according to phone network data [schlapfer2014scaling], which is far from the $N^{2}$ null prediction for a completely well-mixed group.
@@
Furthermore, superlinear exponents, such as the nearly $2$ scaling exponent observed in bacteria, impose a fundamental constraint on system size. With superlinear exponents, there is a maximum size beyond which all components would be regulatory. For bacteria to grow larger, they must fundamentally transform their organizational structure to one that leads to a lower scaling exponent. While the $\approx 2$ exponent is inefficient for scaling up, it also has significant benefits. New genes can be easily added to the bacterial genome along with new regulation in a plug-and-play manner that provides remarkable flexibility to adapt to novel and changing environments. This insight may be transferred to small human organizations, such as start-up companies and local communities. These small organizations are similar to bacteria in the sense that new functions can be easily added with the caveat that everyone interacts directly with everyone else, leading both to a high degree of flexibility and unexpected conflicts. Our theory predicts this form of organization is limited by a predictable critical size. While this study did not collect data on start-up companies, it would be valuable for future research to analyze the regulatory costs and structure of start-up companies on a large scale and examine their transition to a more modular configuration.
@@
Our work makes several contributions to the literature. While most management studies treat regulation as an independent variable, examining its effects on other factors, our research offers a distinct perspective by exploring what determines the level of regulation itself. We conduct a comprehensive empirical comparison across socioeconomic and biological systems, analyzing these systems based on their underlying mechanisms. For the first time, we compare regulatory costs across both biological and socioeconomic systems, identifying their commonalities and differences. Additionally, we introduce a conceptual framework that, despite lacking system-specific details, provides a foundation for understanding baseline regulatory requirements through an optimization process. We hope our work paves the way for a unified science of regulatory costs.
@@ Acknowledgements
This research was supported by the National Science Foundation Grant Award Number EF–2133863. The authors thank Seoul Lee for helping collect and analyze South Korean company data.
@@
Data and methods
Data
Bacteria and Unicellular eukaryotes. Data on genome length and the number of regulatory genes are from van Nimwegen, E. Scaling laws in the functional content of genomes. In Power Laws, Scale-Free Networks and Genome Biology, 236–253 (Springer, 2006)
Cities. Data for occupations in cities were obtained from the Bureau of Labor Statistics in 2017. Data were downloaded from https://www.bls.gov/oes/tables.htm in May 2019. Lawyers are defined according to the 2010 Standard Occupational Classification (SOC) system. The boundaries of urban areas in this dataset use the Metropolitan Statistical Area established by the US Office of Management and Budget (OMB).
Federal agencies. Data on US Federal agencies were obtained from FedScope in 2018. Downloaded from https://www.fedscope.opm.gov/employment_access.asp in Oct 2019. The data provides individual-level information on all US federal government employees, including the federal agency employed under, and the occupation classification. The dataset contains 125 cabinet-level departments and independent agencies. The smallest agencies in this dataset contain two employees, which are Commission for the Preservation of America’s Heritage Abroad and the Northern Border Regional Commission. The largest agencies contain hundreds of thousands of employees—the largest two being the Department of Veterans Affairs, with 391,187 employees, and the Department of the Army, with 249,074 employees. The FedScope data is compiled through human resource software used by the US federal government. Note that not all federal agencies are reported in this data set; however, all agencies available through FedScope are used in our analysis. Supervisors are defined as individuals coded to have supervisor status in the FedScope dataset. The occupation classification in this dataset is according to the OPM’s Handbook of Occupational Groups and Families, available at https://www.opm.gov/policy-data-oversight/classification-qualifications/classifying-general-schedule-positions/occupationalhandbook.pdf. The occupation categories used in our analysis is on the 4-digit level, the finest level in the dataset.
Norwegian Companies The data on Norwegian companies were purchased from Statistics Norway in May 2022. The data are from the year 2019. For confidentiality, the company-level data was aggregated by Statistics Norway according to the following procedure. The companies are first sorted by the number of employees from large to small. For every five companies, both the number of employees and the number of distinct functions are aggregated by taking the mean of the log-transformed variables, i.e., $\text{E}[\log y_{i}]$ where $y_{i}$ is a variable for company $i$ in that size bin. This data only includes companies with five or more employees for confidentiality reasons. Managers are classified according to the International Standard Classification of Occupations (ISCO-08), which is detailed at https://www.ssb.no/en/klass/klassifikasjoner/7/versjon/33. Note that this dataset only includes resident employees in Norway, and it does not include subsidiary companies in countries other than Norway. We filter for a minimum of 50 employees in company size.
Universities. The data on US universities are obtained from the Integrated Postsecondary Education Data System (IPEDS). The data was accessed from https://nces.ed.gov/ipeds/use-the-data. The data used in our analysis are from the year 2016. The type of universities are classified based on the Carnegie 15 classification system, also available in this dataset. Bachelor-level and above institutions include universities classified as ”Baccalaureate,” ”Master,” or ”Doctoral” universities in the Carnegie 15 classification system. Liberal Arts universities are coded as ”Baccalaureate Colleges–Liberal Arts,” Associate Colleges are coded as ”Associates Colleges,” Doctoral universities are coded as ” Doctoral/Research Universities–Extensive” and ”Doctoral/Research Universities–Intensive” in Carnegie 15. In our analysis, we filter for universities with a minimum of 50 employees.
Industries The data on US industries are obtained from the Bureau of Labor Statistics in 2018, downloaded from https://www.bls.gov/oes/tables.htm. Managers and other occupations are defined according to the 2010 Standard Occupational Classification (SOC) system.
Estimation of scaling exponents
The scaling exponents are derived from doing a linear fit after taking the log of both axes. For two variables $x$ and $y$, the exponent $\beta$ and its confidence interval are obtained by regression: $\log y=\beta\log x+\log c$.

### v2

@@ 2.1 Model for baseline expectations of regulatory costs across systems
This prediction of linear scaling is supported by the observation that in unicellular eukaryotes, the number of regulatory genes, is indistinguishable from linear [van2003scaling]. Unlike bacteria, unicellular eukaryotes have various internal spatial partitions, and the partitioning of genes between the nucleus and mitochondria along with the separation of genes into different chromosomes and with a complex system of regulation that is still being uncovered [fritz2019chromosome, li2018gene, cremer2010chromosome]. The transition from prokaryotes to eukaryotes illustrates how the internal structure of organisms, hierarchy, and partitioning can alter the requirements for regulation.
Considering the transitions across systems of different sizes, and that the cost per individual likely varies with system size, our model suggests that the number of regulators generalizes to
@@
where $r_{0}$ is the minimum cost for the smallest size ($N=1$), and $\beta$ is the scaling exponent. This result is in keeping with the general observation that other metrics in these systems not directly related to regulation follow such power-law behavior with non-trivial exponents. Our model presented above provides base-line expectations for the values of $\beta$, linking its value to system organization: $\beta=2$ for a compartment-free, well-mixed system, and $\beta=1$ for a system with both regulators and compartments. Consequently, we can interpret observed exponents that are closer to $2$ as being well-mixed and compartment-free. Exponents that are close to $1$ indicate a system optimized on compartments with a mixed strategy. An exponent significantly larger than $1$ and significantly smaller than $2$ indicates the transition from regulators only to some compartmentalization. Exponents significantly less than $1$ indicate the regime where compartments are expanding and regulators are becoming cumbersome and saturating, or where there are certain economies of scale in the unit costs.
@@ 2.2 Empirical results
Figure 3: Examples of scaling of regulatory functions in bacterial genomes, cities, and Norwegian companies (a) The number of transcription regulatory genes vs. the number of genes in the genome of bacterial cells. (b) The number of lawyers vs. working population in US Metropolitan Statistical Areas. (c) The number of managers in companies in Norway.
We collect data on regulatory components and system size in both biological and socioeconomic systems to examine how regulatory costs scale with size. In biological systems, we focus on bacteria and unicellular eukaryotes, specifically collecting data on the number of transcription regulatory genes in their genomes. Transcription regulatory genes are those that determine the timing and environmental conditions for gene expression by producing proteins that bind to other genes. We use data reported by [van2003scaling] for these biological systems (see Supplemental Information for a comprehensive discussion). For socioeconomic systems, we gathered data on the number of lawyers in cities and the number of managers in Norwegian and Korean companies, US federal government agencies, and various types of US universities [youn2016scaling, gulati2023]. We also synthesize results from another study measuring the number of administrators for Wikipedia pages [yoon2023makes]. For human systems, system size is measured by the population of the entity, such as the number of employees for companies and universities, and the number of editors for Wikipedia pages. The scaling exponents of these systems are estimated using Eq. 5, with $\beta$ representing the scaling exponent. For detailed information on data sources and statistical methods used for estimation, see Supplementary Information.
@@
In the data we gathered, regulatory costs scale with system size across diverse systems following similar power law behavior remarkably closely. Three examples are shown in Fig. 3: regulatory genes in bacteria, lawyers in cities, and managers in Norwegian companies. As predicted, the scaling exponents vary across system types, with regulatory genes in small bacteria having the highest exponent at $1.65$. The number of lawyers scales superlinearly with the urban population, with an exponent of $1.28$ (Fig. 3B). This is similar to prokaryotic regulatory genes but with a lower scaling exponent, likely due to modularization in the interaction network structure of cities. Theories of urban scaling based on these constraints have successfully predicted similar exponents for many socioeconomic outputs driven by interactions [bettencourt2013origins, yang2019modeling]. In contrast to cities, other human organizations—such as government agencies, companies, and universities—typically exhibit a high degree of hierarchical structure, leading us to expect different scaling exponents closer to unity. In these organizations, managers play a crucial role in coordinating efforts and mitigating conflicts among subordinates. Thus, the scaling exponent for managers in Norwegian companies is 0.91 in good agreement with our prediction (Fig. 3C).
@@
The scaling exponents for all data we have gathered, including the number of managers in all sectors of US universities, Norwegian and Korean companies, and US federal government agencies, as well as those in biological systems, are summarized in Table 1. In all the modular systems, the number of managers scales sublinearly with small variations—from $0.94$ for Federal agencies (highest exponent) to $0.72$ for Korean companies (lowest exponent). The sublinear scaling shown in the data suggests the span of control, the number of subordinates per manager, increases with organization size, which aligns with previous studies in management science [boyne2013burdened]. This finding is contrary to the popular belief that larger organizations are less efficient in coordination costs measured in managers per capita [williamson1967hierarchical]; in fact, the data suggests they are more efficient at exploiting economies of scale. Even in Wikipedia, a supposedly decentralized system, the interaction network among editors has been shown to be highly modularized [yoon2023makes]. Interactions tend to cluster around specific topics of interest, allowing a smaller number of administrators to effectively manage issues that arise. As such, despite the decentralized nature of the system, administrators naturally oversee specific modules and thus scale sublinearly with the total number of contributors.
@@
Table 1: The scaling exponents ($\beta$) and 95% confidence intervals for regulatory functions in various systems. With more well-mixed internal structures, such as bacterial cells and cities, regulatory functions tend to be superlinear. With more modular internal structures, such as companies and universities, the scaling exponents are linear to sublinear.
@@ 2.3 Function diversity associated with scaling deviations
Figure 4: Data for Industries and Federal agencies colored by occupations’ normalized entropy, a metric for the diversity of functions in the system. Higher values denote more diversity. The inset shows the correlation between the normalized entropy and scaling deviation. Entities that are more diverse have more managers than would be expected of their size. The Pearson correlation is 0.59 for industries and 0.20 for federal agencies.
Motivated by our theory, we predict that greater functional diversity is associated with higher regulatory costs. To test this prediction, we quantify functional diversity across a range of systems and demonstrate that it is positively associated with the scaling residuals. In other words, systems with higher functional diversity exceed the expected number of regulators as would predicted by the scaling curve, while those with lower functional diversity fall short.
We quantify functional diversity in a system by analyzing the distribution of occupations within it. For this analysis, we utilize individual-level occupation information for US federal government agencies. To ensure robustness, we also test our predictions on companies. Although occupation information is not available to us at the company level, we study companies aggregated into industries, for which we have detailed accounts of the distribution of occupations. We measure functions using the finest occupational categories available in these datasets. For more details on data sources and occupation definitions, see Supplementary Information.
@@
We also compute the scaling residual [bettencourt2010urban] for each system, which quantifies the extent to which a system over- or under-performs relative to the scaling curve. The scaling residual for entity $i$, $\xi_{i}$, is defined as, $\xi_{i}=\log r_{i}-\log r(N_{i})$, where $r_{i}$ is the number of regulators for entity $i$ in the data, and $r(N_{i})$ is the regulators expected of its size according to Eq. 5.
Figure 4 illustrates the relationship between managers and employees, with each entity colored according to its normalized entropy, where higher values indicate greater function diversity [yang2022scalinguniversalityfunctiondiversity]. For both federal agencies and industries, entities with higher function diversity tend to have more managers than expected for their size. The insets of Fig. 4 display the correlation between scaling residuals and normalized entropy. The correlation is 0.59 for industries ($p<0.001$), and 0.20 for federal agencies ($p=0.030$). These results support our initial hypothesis that greater function diversity is associated with greater regulatory costs as a result of greater coordination requirements.
@@ 3 Discussion
We have proposed a conceptual framework for unifying the conceptualization of regulatory costs across biological and socioeconomic systems by examining the interactions of fundamental features, size and internal structure. We also conducted an empirical analysis of regulatory costs across these systems. Our findings indicate that the variation of regulatory functions with system size depends significantly on the system’s structure. The exponents range from 1.65 for regulatory genes in bacteria, which have a well-mixed internal structure and genes located on a single chromosome, to sublinear scaling for managers in hierarchically organized human organizations, such as federal agencies and universities. By characterizing regulatory functions based on their underlying mechanisms rather than the system type, our work represents a first step toward a unified understanding of regulatory costs across diverse systems.
@@
Our conceptual model is based on the simplest assumptions about cost factors, intended to lay the groundwork for more advanced theories in the future. These future models should incorporate more accurate cost estimates tailored to different system types through empirical measurements. For example, in hierarchical human organizations, the cost of regulators within compartments should also be factored into future models. In biological systems, more precise cost calculations should consider the frequency of gene expression, which plays a critical role in determining the energetic cost in the cell [kempes2016evolutionary, kempes2017drivers, lynch2015bioenergetic].
@@
It is critical to note that the greater scaling exponents of biological systems do not imply that biological systems are less efficient than socioeconomic ones. Our framework shows that for small systems, compartments are too expensive, and it is better to simply add regulators in a superlinear fashion. Compartmentalization and structure in biology is expensive—intricate physical structures need to be developed and maintained and this is only beneficial after a critical size. Extending this idea, complex physical infrastructure also needs to be developed to enable organs in animals, where genes only get expressed in a certain tissue. While bacteria need to carry a rapidly increasing number of regulatory genes with increasing size, they save on the energetic cost of creating compartmentalization within the cell. In comparison, developing structure in social systems does not necessarily incur a physical cost. For example, a company’s CEO can decide to create a new division in the company without employing new physical separations between divisions, and the re-organization can be accomplished in a matter of weeks. Humans are also naturally creatures of groups with limited social capacity, making modularity a common characteristic in many social systems, even decentralized ones such as cities. Indeed, the social interactions scale with $N^{1.2}$ according to phone network data [schlapfer2014scaling], which is far from the $N^{2}$ null prediction for a completely well-mixed group.
@@
Furthermore, superlinear exponents, such as observed in bacteria, impose a fundamental constraint on system size. With superlinear exponents, there is a maximum size beyond which all components would be regulatory. For organisms to grow beyond this limit, they must fundamentally transform their organizational structure to one that leads to a lower scaling exponent. While the superlinear exponent is inefficient for scaling up, it also has significant benefits. New genes can be easily added to the bacterial genome along with new regulation in a plug-and-play manner that provides remarkable flexibility to adapt to novel and changing environments. This insight may be transferred to small human organizations, such as start-up companies and local communities. These small organizations are similar to bacteria in the sense that new functions can be easily added with the caveat that everyone interacts directly with everyone else, leading both to a high degree of flexibility and unexpected conflicts. Our theory predicts this form of organization is limited by a predictable critical size. While this study did not collect data on start-up companies, it would be valuable for future research to analyze the regulatory costs and structure of start-up companies on a large scale and examine their transition to a more modular configuration.
@@
Our work provides a unified, first-principles framework for understanding regulatory costs across biological and socioeconomic systems. In contrast to most studies in management and organizational science, which rely on regression-based analyses of correlates to regulatory burden, we develop an optimization-based model grounded in fundamental mechanisms by which systems manage adverse interactions: toleration, regulation, or compartmentalization. This model yields baseline expectations for regulatory costs as a function of system size and internal structure.
A central innovation is our cross-system comparison, linking biological and human systems through shared structural constraints. While prior work has studied regulation within each domain separately, we show for the first time that regulatory functions—from regulatory genes in genomes to managers in organizations—follow systematic scaling relationships with size, shaped by their internal architecture. This cross-system comparison can bring insights from biological systems to help understand social systems. In human organizations, regulatory costs have been rising, yet it is unclear whether such costs are an unavoidable consequence of organizational scale and complexity or the result of mismanagement and redundancy. By drawing on biological systems, which have been shaped by evolutionary optimization and where necessary regulation is more clearly defined, we uncover general tradeoffs that apply across systems, and can be used to understand essential versus excessive in human organizations.
We also show that deviations from expected scaling can be predicted by functional diversity, indicating that interaction complexity drives additional regulatory needs. Together, these contributions move beyond domain-specific explanations and toward a general, mechanistic science of regulation across complex systems.
@@ Acknowledgements
This research was supported by the National Science Foundation Grant Award Number EF–2133863. The authors thank Seoul Lee for helping collect and analyze South Korean company data. H.Y. acknowledges the NRF Global Humanities and Social Sciences Convergence Research Program (2024S1A5C3A02042671) and the support from the Institute of Management Research at Seoul National University.
Code and Data Availability
The data used in this study and code used to analyze the data are included as a replication file with this submission. The Supplementary Information details method for data collection.